OpenStudy (samara8954):

Sam opened a money-market account that pays 2% simple interest. He started the account with $7,000 and made no further deposits. When he closed the account, he had earned $420 in interest. How long did he keep his account open? 3 years 6 years 4 years 5 years

OpenStudy (samara8954):

@cwrw238

OpenStudy (samara8954):

help I will medal @chefducky

OpenStudy (anonymous):

i dunno

OpenStudy (samara8954):

@SyedMohammed98

OpenStudy (samara8954):

@natsent06

OpenStudy (dtan5457):

For one year interest you use 0.02 times 7000=$140 1 year=140 interest how many for 420 interest??

OpenStudy (samara8954):

I don't know help @dtan5457

OpenStudy (dtan5457):

420 divided by 140 is..?

OpenStudy (samara8954):

3

syed98 (syedmohammed98):

its 3 here the solution ar(n)=total a is the first number which is 7000 r is the common ratio which is 100%+2%=102 102 divided by 100 giving u 1.02(interest) so ar(n)=sum n is written as power.....and is the number of years so 7000*1.02(n)=(7000+420) 1.02(n)=7420/7000 1.02(n)=1.06 log10 1.02(n)=log10(1.06) ( multiply log10 with both) n=log10(1.06)/log10 (1.02) u get 2.94 which is round off to 3 so 3 years.. hope u get it :)

OpenStudy (dtan5457):

HAHA I DIED ^

syed98 (syedmohammed98):

lol